### v1

# Test-Time Scaling of Reasoning Models for Machine Translation

Abstract
Test-time scaling (TTS) has enhanced the performance of Reasoning Models (RMs) on various tasks such as math and coding, yet its efficacy in machine translation (MT) remains underexplored. This paper investigates whether increased inference-time computation improves translation quality. We evaluate 12 RMs across a diverse suite of MT benchmarks spanning multiple domains, examining three scenarios: direct translation, forced-reasoning extrapolation, and post-editing. Our findings show that for general-purpose RMs, TTS provides limited and inconsistent benefits for direct translation, with performance quickly plateauing. However, the effectiveness of TTS is unlocked by domain-specific fine-tuning, which aligns a model’s reasoning process with task requirements, leading to consistent improvements up to an optimal, self-determined reasoning depth. We also find that forcing a model to reason beyond its natural stopping point consistently degrades translation quality. In contrast, TTS proves highly effective in a post-editing context, reliably turning self-correction into a beneficial process. These results indicate that the value of inference-time computation in MT lies not in enhancing single-pass translation with general models, but in targeted applications like multi-step, self-correction workflows and in conjunction with task-specialized models.
Test-Time Scaling of Reasoning Models for Machine Translation
Zihao Li,1 Shaoxiong Ji,2,3 Jörg Tiedemann1 1University of Helsinki  2 University of Turku  3 ELLIS Institute Finland firstname.lastname@{1helsinki.fi, 2utu.fi}
1 Introduction
Figure 1: Illustration of the effectiveness of test-time scaling in reasoning models for machine translation. (1) TTS for general-purpose RMs yields only a small initial performance gain, but quickly plateauing as increased inference cost. (2) Forcing RMs to reason beyond their natural stopping point degrades quality by introducing noise. (3) In contrast, TTS becomes effective when applied to RMs specifically developed for MT. (4) TTS shows improvements in post-editing workflows. All these highlight TTS’s value in MT lies in task-specialized models and multi-step self-correction, rather than as a robust strategy for enhancing single-pass translation with general-purpose RMs.
Large language models (LLMs) have dramatically advanced machine translation (MT), evolving from statistical and neural paradigms to systems capable of handling diverse languages, domains, and complexities with unprecedented accuracy (Lyu et al., 2023; Kocmi et al., 2024; Zhu et al., 2024; Cui et al., 2025). Recent developments in Reasoning Models (RMs)—models designed to incorporate structured reasoning processes like Chain-of-Thought (CoT)—have further transformed MT by reframing it as a cognitive task requiring contextual analysis, cultural adaptation, and self-reflection (Liu et al., 2025). For instance, RMs can resolve ambiguities in stylized texts and maintain coherence across documents, thereby outperforming traditional LLMs in semantically demanding scenarios (Ye et al., 2025).
Test-time scaling (TTS) has emerged as a transformative approach for enhancing model performance, which allocates additional computational resources during inference to enhance performance without requiring model retraining or parameter expansion (Snell et al., 2024). The effectiveness of TTS has been particularly pronounced for RMs such as DeepSeek-R1 (Guo et al., 2025), Gemini 2.5 (Comanici et al., 2025), and OpenAI’s o-Series (Jaech et al., 2024; OpenAI, 2025), which have achieved breakthrough performance on challenging benchmarks by extending their reasoning chains. Moreover, relatively small RMs have demonstrated impressive results on mathematical and coding tasks through TTS (Muennighoff et al., 2025; Li et al., 2025), suggesting that inference-time computation can partially compensate for limited model capacity.
Nevertheless, applying TTS to RMs for MT introduces distinct challenges and untapped potential that warrant deeper exploration. Unlike math or coding tasks, where correctness can often be objectively determined, MT demands not only linguistic accuracy but also reasoning over cultural nuances, domain-specific terminology, and long-range dependencies, areas where unstructured compute scaling may yield diminishing returns. Moreover, interventions like forced extrapolation (e.g., inserting “wait" tokens to extend reasoning) could disrupt natural deliberation, potentially introducing noise. In post-editing (PE) contexts, where models refine their own drafts, TTS might unlock iterative improvements, though this demands rigorous testing across varied benchmarks.
This paper investigates these open questions through three research questions:
  • RQ1: How effective is test-time scaling for MT? We examine whether increased inference computation reliably boosts translation quality across general-purpose and fine-tuned MT-specific RMs.
  • RQ2: Does extrapolation by inserting “wait” forcibly help? We investigate if overriding models’ natural stopping points, which further scales up the inference computation, enhances or hinders performance.
  • RQ3: Does test-time scaling work in post-editing? We evaluate TTS in self-correction scenarios, assessing its role in refining initial translations.
To address these questions, we assemble a comprehensive array of MT benchmarks encompassing literary, biomedical, cultural, commonsense, constrained terminology, and retrieval-augmented domains. We assess 12 RMs, spanning open-source series (Qwen-3 (Yang et al., 2025), Cogito (Deep Cogito, 2025), DRT (Wang et al., 2025a)) and the proprietary Grok-3-Mini.
Our key contributions and findings are as follows:
  • We demonstrate that for general-purpose RM, TTS provides limited and inconsistent benefits for direct machine translation. After small initial improvements at very low budgets, performance plateaus across metrics and datasets, indicating that “more thinking” alone is not a robust path to better translations.
  • We show that the effectiveness of TTS is unlocked by domain-specific fine-tuning, which aligns the model’s reasoning process with task requirements. For DRT models fine-tuned on specific domain data, performance improves with budget on in-domain tasks and saturates once models naturally stop increasing their internal token usage, suggesting an emergent alignment between optimal reasoning depth and task demands. This alignment largely disappears out of domain.
  • We find that forcing a model to reason by inserting a single “wait” beyond its natural stopping point consistently degrades translation quality, highlighting the importance of the model’s intrinsic deliberation process.
  • We establish that TTS is highly effective in a post-editing context, in which the inference cost is higher than the cost of direct translation. TTS turns self-correction into a reliably beneficial process.
These findings provide implications for deploying RMs in production MT systems and highlight the critical interplay between model capacity, task-specific training, and inference-time computation in determining when and how test-time scaling benefits translation quality.
2 Experimental Setup
2.1 Datasets
To comprehensively evaluate the reasoning capabilities and scaling properties of models at test time, we curated a diverse suite of eight machine translation benchmarks. These datasets span multiple domains, granularities, and languages, targeting a wide spectrum of reasoning challenges (Table 1).
For tasks requiring deep contextual and stylistic understanding, we use three literary benchmarks: the document-level WMT24-Literary (Wang et al., 2024b), paragraph-level LitEval-Corpus (Zhang et al., 2025), and sentence-level MetaphorTrans (Wang et al., 2025a). These datasets are rich in complex linguistic phenomena, cultural references, and figurative language (similes and metaphors), demanding sophisticated reasoning to preserve literary style and meaning. Similarly, the WMT23/24-Biomedical (Neves et al., 2023, 2024) benchmark tests reasoning within a specialized domain, demanding accurate translation of technical terminology from PubMed abstracts.
To probe more targeted reasoning abilities, we incorporate four specialized datasets. CAMT (Yao et al., 2024) assesses cross-cultural reasoning on expressions requiring cultural adaptation. Commonsense-MT (He et al., 2020) comprises subsets targeting lexical, contextless syntactic, and contextual syntactic ambiguities, each requiring commonsense reasoning. RTT (Zhang et al., 2023) evaluates constrained reasoning by requiring models to correctly translate specific terminology under highly constrained conditions. Lastly, RAGTrans (Wang et al., 2024a) examines a model’s capacity to reason over and integrate retrieved external evidence into its translation. Collectively, these benchmarks provide a rigorous and multifaceted framework for analyzing the effects of scaling on translation reasoning.
Domain Datasets Granularity Languages Language Pair(s) Sample Size
Literature WMT24-Literary (Wang et al., 2024b) Document-level ZH, DE, RU 3 43
MetaphorTrans (Wang et al., 2025a) Sentence-level ZH, EN 1 2000
LitEval-Corpus (Zhang et al., 2025) Paragraph-level ZH, EN, DE 4 187
Biomedical WMT24-Biomedical (Neves et al., 2024) Document-level EN, DE, ES, FR, IT, PT, RU 12 600
WMT23-Biomedical (Neves et al., 2023) Document-level 585
Culture CAMT (Yao et al., 2024) Sentence-level EN, ES, FR, HI, TA, TE, ZH 7 6948
Commonsense Commonsense-MT (He et al., 2020) (Lexical Ambiguity) Sentence-level ZH, EN 1 400
Commonsense-MT (Contextless Syntactic Ambiguity) Sentence-level 450
Commonsense-MT (Contextual Syntactic Ambiguity) Sentence-level 350
Terminology RTT (Zhang et al., 2023) Sentence-level EN, DE 2 100
Misc. RAGTrans (Wang et al., 2024a) Sentence-level ZH, EN 1 1999
Table 1: Overview of the MT benchmarks used in our evaluation.
2.2 Models
Our evaluation encompasses 12 RMs, including 11 open-source models from three distinct families and one proprietary model for comparison. The open-source models investigated are as follows:
  • Qwen-3: Six models from this family were selected, with parameter sizes of 0.6B, 1.7B, 4B, 8B, 14B, and 32B (Yang et al., 2025). These models are hybrid reasoning LLMs that support seamless switching between a standard generation mode and a deliberative reasoning mode.
  • Cogito: Two models, sized 3B and 8B, were included (Deep Cogito, 2025). Cogito-3B and Cogito-8B are trained on top of Llama-3.2-3B and Llama-3.1-8B (Dubey et al., 2024), respectively, and similarly implement hybrid reasoning capabilities with controllable switching between generation and reasoning modes.
  • DRT: Three models from this family were evaluated. These models are fine-tuned from existing LLMs using the training set of MetaphorTrans (Wang et al., 2025a). Specifically, DRT-7B, DRT-8B, and DRT-14B are built upon Qwen2.5-7B-Instruct (Qwen et al., 2024), Llama-3.1-8B-Instruct, and Qwen2.5-14B-Instruct respectively.
In addition to the open-source models, we included the proprietary model Grok-3-Mini. This model provides a tunable reasoning_effort parameter (equivalent to reasoning budget but can only be set to low or high) to control the amount of deliberation performed prior to generating a response.
2.3 Evaluation Metrics
We assess translation quality using a suite of automatic metrics, encompassing both reference-based and reference-free approaches, alongside a specialized LLM-based judge for literary texts.
COMET Metrics.
For a standardized assessment, we employ two variants from the COMET framework (Rei et al., 2020): the reference-based COMET-22 (Rei et al., 2022a) and the reference-free COMETKiwi-22 (Rei et al., 2022b).
LLM as Judge.
[…]
2.4 Budget Forcing
We regulate test-time reasoning with a logits processor that enforces a thinking-token budget inside a <think>…</think> span. While in this span, the processor counts tokens, softly encourages closure near 95% of the budget by upweighting newline and </think>, then deterministically emits a newline (penultimate step) and </think> (final step) at the budget limit before continuing normal answer decoding.
Conversely, to probe extrapolation, we optionally insert a single “wait” token if the model attempts to stop: specifically, after at least 5 thinking tokens, if </think> is the next token from the output of the argmax function and the budget is not yet exhausted, we override the next token to “wait” once and resume unconstrained decoding. We insert “wait” at most once.
2.5 Post-Editing
Post-editing (or self-correction) involves two-stage translation, i.e., stage one of direct translation and stage two that corrects or post-edits the direct translation, which enables models to review and refine their own outputs (Feng et al., 2025; Wang et al., 2024c; Li et al., 2024). We explore two prompting strategies to guide this self-correction process, with full details provided in Appendix D. The first is a standard PE prompt, which we term “No QS” (No Quality Score). It provides the model with only the source text and its own draft translation to be refined. The second is an enhanced prompt, “QS” (with Quality Score), which additionally includes a numerical quality score of the draft, calculated as the average of the GRB and GRF scores from the initial translation. This provides the model with an explicit signal about the quality of the translation it needs to correct, potentially guiding a more targeted reasoning process.
Figure 2: Average GRB scores of Qwen-3 and Cogito models across all datasets with varying thinking budgets.
3 Results and Analysis
3.1 Effectiveness of Test-Time Scaling
General-Purpose Models Show Limited Gains from Increased Budget.
Our initial investigation focused on the efficacy of test-time scaling for general-purpose RMs, including the Qwen-3 and Cogito families, as well as the proprietary Grok-3-Mini model. These models were evaluated “out-of-the-box" without any fine-tuning on our benchmark datasets. Figure 2 plots the average GRB scores across all datasets against varying thinking-token budgets for the Qwen-3 and Cogito model series. After a small initial performance gain when moving from a zero budget to a minimal budget (e.g., 100 tokens), the models’ performance curves almost completely plateau in most cases. Further increases in the budget up to 2000 tokens yield no meaningful improvements, indicating that simply allocating more computational steps does not enable the models to produce more refined or accurate translations when they lack specific task-related knowledge.
This conclusion is further corroborated and nuanced by our analysis of the Grok-3-Mini model. We analyzed its performance using both reference-based (GRB) and reference-free (GRF) metrics, visualized in Figure 3(a) and Figure 3(b), respectively. Both metrics reveal a highly inconsistent, dataset-dependent impact. For instance, while higher effort improves scores on CommonsenseMT-Lexical across both GRB (+0.450) and GRF (+0.376), it significantly degrades performance on CommonsenseMT-Contextless in both cases (-0.780 for GRB, -0.367 for GRF). Crucially, the average effect across all datasets is negligible and even flips its sign depending on the metric: the mean GRB delta is a slightly negative -0.064, while the mean GRF delta is a slightly positive +0.033. Given that both scores are on a 100-point scale, these near-zero average changes underscore that the potential benefits and drawbacks of increased computational effort effectively cancel each other out, leading to no reliable overall improvement.
Therefore, our analyses of both open-source models with varying budgets and a proprietary model with different effort levels converge on a single conclusion: for general-purpose LLMs without specific in-domain training, test-time scaling is not a robust strategy for enhancing machine translation performance.
Figure 3: Performance of Grok-3-mini across tasks, showing the difference between high- and low-effort reasoning. Subfigure (a) reports results under the GRB metric, and (b) shows results under GRF.
In-Domain Fine-Tuning Unlocks the Benefit of Test-Time Scaling.
In contrast to the general-purpose models, the DRT models reveal that the effectiveness of test-time scaling is highly contingent on domain-specific training, which appears to create an efficient alignment between reasoning effort and performance. These models were fine-tuned on the training set of MetaphorTrans (in-domain), while LitEval-Corpus and WMT24-Literary serve as related but out-of-domain literary benchmarks. Figure 4 visualizes the translation quality (GEA score, right y-axis) and the actual number of thinking tokens generated by the models (left y-axis).
On the in-domain MetaphorTrans task, we observe a clear and consistent positive correlation between the thinking budget and translation performance. As the thinking budget increases from 100 to around 500 tokens, both the number of generated tokens and the GEA scores steadily rise.
However, beyond a 500-token budget, a critical pattern emerges: the models stop generating more thinking tokens, and concurrently, their performance plateaus. We hypothesize that the fine-tuning successfully aligns the model’s reasoning behavior with the effective reasoning boundary of the task.
This efficient alignment vanishes on the other out-of-domain literary translation tasks. The most striking counterexample is the document-level WMT24-Literary task. Here, the actual thinking tokens continue to scale almost linearly with the budget, indicating the models are using the provided extra capacity to “think" longer. Yet, this extended reasoning does not translate into better performance; the GEA scores remain erratic and show no consistent improvement. This disconnect suggests the models are engaged in unproductive or unfocused reasoning, “spinning their wheels" without the specialized knowledge required for this different type of literary translation. This dichotomy underscores our central argument: test-time scaling is most effective when fine-tuning has equipped a model with not only domain-specific knowledge but also an efficient strategy for how and when to apply it.
Figure 4: Performance (dashed lines, right axis) and actual generated thinking tokens (solid lines, left axis) of DRT models across 3 literary translation tasks.
3.2 Forced Extrapolation of Reasoning Degrades Performance
Building on the observation that models possess a natural reasoning length, we next address RQ2: Does performance improve if we force a model to think longer? We test this by applying the “wait” token extrapolation method, detailed in Section 2.4, to prompt continued reasoning when the models are about to stop the thinking process.
The results, averaged across all datasets for the Qwen-3 and Cogito models, are presented in Table 2. The findings are unambiguous. First, as evidenced by the “Thinking Length” columns, the intervention was effective in its primary goal: it consistently and significantly extended the models’ reasoning chains, often by over 100-200 tokens.
However, this artificially prolonged reasoning process did not translate into better translations. In fact, it was overwhelmingly detrimental. Across all four metrics (COMET, COMETKiwi, GRB, and GRF), the “wait” token intervention consistently reduces performance: under both the 1000- and 2000-token budgets, 55 of the 64 metric scores across eight models dropped after forced extrapolation. While there are a few isolated instances of negligible score increases in one metric (e.g., Qwen3-4B on GRB with a 2000-token budget), these are exceptions that are contradicted by decreases in other metrics for the same model.
This leads to a clear conclusion: a model’s decision to terminate its reasoning chain is a meaningful signal. It indicates that the model has reached what it considers to be a sufficient state of deliberation for the given task. Forcing the model to continue to reason beyond its own stopping point appears to introduce noise, repetition, or less relevant reasoning steps, which ultimately harms the quality of the final translation. In short, we find that forced extrapolation is a counterproductive strategy for improving translation quality.
Model Budget Thinking Length COMET COMETKiwi GRB GRF
Before After Before After Before After Before After Before After
Qwen3-0.6B 1000 426 519 0.6959 0.6904 0.6155 0.6026 68.8577 68.6333 67.2944 67.1203
[…]
Qwen3-1.7B 1000 655 820 0.7687 0.7475 0.6851 0.6543 83.6481 83.5185 83.1170 82.7297
[…]
Qwen3-4B 1000 737 873 0.7914 0.7738 0.7092 0.6835 89.7781 89.5598 89.7522 89.5203
[…]
Qwen3-8B 1000 741 878 0.7979 0.7865 0.7180 0.6954 91.6706 91.7601 91.8085 91.7027
[…]
Qwen3-14B 1000 668 812 0.7992 0.7924 0.7184 0.7022 92.5609 92.5677 92.5242 92.5148
[…]
Qwen3-32B 1000 700 835 0.8025 0.7828 0.7233 0.6957 92.6026 92.5537 92.6328 92.5322
[…]
Cogito-3B 1000 577 695 0.7071 0.7063 0.6313 0.6310 84.0098 83.8203 83.1502 82.5440
[…]
Cogito-8B 1000 568 697 0.7678 0.7658 0.6824 0.6807 89.6573 89.3849 89.2655 88.8525
[…]
Table 2: Effect of forcibly inserting a “wait” token to extend the reasoning process. The “Before” columns show standard generation, while “After” shows results from the intervention.
3.3 Test-Time Scaling is Effective for Post-Editing
Finally, we investigate RQ3 by evaluating the effectiveness of test-time scaling in a self-correction post-editing scenario. For this task, we define the baseline as the translation generated by each Qwen-3 model with a zero thinking budget in our prior experiments. Subsequently, we task the same model with refining its own translation, applying thinking budgets of 0, 500, and 1000 tokens. The detailed results are presented in Table 12 and Table 13, with trends visualized in Figure 5.
In a striking contrast to its ineffectiveness in direct translation, test-time scaling proves to be a highly effective strategy for post-editing, reliably elevating translation quality above the original baseline for most models. The effect is most pronounced for mid-sized models, as shown in both the GRB (Figure 5(a)) and GRF (Figure 5(b)) plots. For models in the 1.7B to 14B parameter range, applying post-editing with a zero budget often yields results that are similar to or worse than the original translation. Increasing the budget to 500 or 1000 tokens consistently pushes performance significantly above this baseline, demonstrating that a thinking budget is crucial for turning self-correction into a reliably beneficial process.
However, this scaling trend does not hold for the extremes of the model family. The smallest model, Qwen3-0.6B, displays erratic behavior, with its performance fluctuating without clear improvement as the budget increases. Conversely, the largest model, Qwen3-32B, already surpasses the baseline with a zero-budget correction, and additional thinking time provides no further gains, suggesting it performs near its peak without extended deliberation.
A comparison of the two prompting strategies further highlights the importance of the thinking budget. At a zero-token budget, the “QS” prompt (with quality score) generally underperforms the “No QS” prompt. However, once the budget is increased to 500 or 1000 tokens, their performances converge and become nearly indistinguishable. This demonstrates that while prompting strategy matters, it is the allocation of a computational budget that is the key factor for post-editing to reliably improve upon the initial translation.
(a) GRB scores for post-editing across Qwen-3 models.
(b) GRF scores for post-editing across Qwen-3 models.
Figure 5: Effectiveness of test-time scaling in post-editing scenario.
4 Related Work
Machine Translation with Large Language Models
LLMs have rapidly advanced machine translation, with growing emphasis on open-weight systems that balance translation quality, multilingual coverage, and general-purpose capabilities. Seed-X (Cheng et al., 2025) introduces a 7B-parameter open-source family of translation-oriented LLMs trained on large-scale monolingual and bilingual corpora across 28 languages, achieving performance competitive with closed-source systems such as GPT-4o (Hurst et al., 2024) and Gemini-2.5 (Comanici et al., 2025). Similarly, Hunyuan-MT (Zheng et al., 2025) develops 7B-parameter models: Hunyuan-MT and Hunyuan-MT-Chimera, the latter integrates multiple outputs under a “slow thinking” paradigm to yield higher-quality translations, and ranks first in the WMT2025 shared task across 30 of 31 directions. Tower+ (Rei et al., 2025) addresses the trade-off between translation specialization and general-purpose ability by combining continued pretraining, supervised fine-tuning, preference optimization, and reinforcement learning with verifiable rewards.
Machine Translation with Reasoning Models
Recent research has explored how RMs can be adapted to MT, particularly in linguistically and culturally challenging domains. Wang et al. (2025a) propose Deep Reasoning Translation (DRT), which leverages long chain-of-thought (CoT) reasoning within a multi-agent framework to tackle similes and metaphors in English–Chinese literary translation. The resulting models surpass standard LLMs by synthesizing long-thought training data. Building on this, DeepTrans (Wang et al., 2025b) applies reinforcement learning (RL) with carefully designed reward functions targeting both translation fidelity and reasoning quality, showing that RL without labeled pairs can significantly boost performance. ExTrans (Wang et al., 2025c) complements this direction with an exemplar-enhanced RL approach that employs a stronger RM (DeepSeek-R1) as a reward reference. ExTrans achieves state-of-the-art results in English–Chinese literary MT, and its multilingual extension (mExTrans) scales effectively to 90 directions with lightweight reward modeling. Beyond literary MT, R1-T1 (He et al., 2025) generalizes reasoning-based MT by modeling six CoT templates inspired by human translator strategies. Through RL, it enables self-evolving reasoning trajectories, improving performance across diverse domains and low-resource languages.
Test-Time Scaling
Recent studies have examined the potential of test-time scaling. Tan et al. (2025) propose a best-of-$N$ reranking framework where multiple translation candidates are generated and the best one is selected using a quality estimation model. They show that smaller models can, through TTS, match or even surpass larger model. For example, a 14B model with $N\!\approx\!8$ achieves parity with a 72B model at $N=1$ while requiring substantially less GPU memory. Beyond MT, Son et al. (2025) analyze TTS in multilingual mathematical reasoning, finding that outcome and process reward modeling, as well as budget forcing, yield notable gains in English but limited improvements across 55 languages, highlighting cross-lingual fragility. Tran et al. (2025) study low-resource reasoning tasks and propose English-pivoted CoT generation, where reasoning occurs in English before producing final answers in the target language, yielding substantial accuracy improvements. Yong et al. (2025) further study cross-lingual reasoning with English-centric RMs, finding that scaling inference budgets with long CoTs improves multilingual mathematical reasoning and even allows smaller models to outperform larger baselines, but also highlighting language-mixing behaviors and weaker generalization to cultural commonsense domains.
5 Conclusion
In this work, we systematically explored the application of test-time scaling (TTS) to reasoning models (RMs) for machine translation (MT), addressing three core research questions through extensive experiments across diverse benchmarks, models, and evaluation metrics.
Our findings reveal that TTS offers limited value for general-purpose RMs in direct translation tasks, where performance quickly plateaus after minimal initial gains, underscoring that additional inference-time computation alone cannot compensate for a lack of task-aligned reasoning strategies. In contrast, domain-specific fine-tuning emerges as a pivotal enabler, allowing TTS to yield consistent improvements on in-domain tasks until models reach their natural reasoning depth, beyond which further scaling provides no benefit. This highlights an emergent efficiency in fine-tuned models, where optimal deliberation aligns with task demands, though such alignment erodes out-of-domain. Furthermore, forcibly extending reasoning via interventions like “wait” tokens consistently degrades quality, emphasizing the importance of respecting a model’s intrinsic stopping points. Finally, TTS proves particularly potent in post-editing scenarios, transforming self-correction into a reliable mechanism for refining initial drafts, especially for mid-sized models when paired with adequate budgets.
The implications of this work are twofold. First, for practitioners, simply allocating more inference compute to general-purpose models is an inefficient path to better translations. Instead, resources are better invested in targeted fine-tuning, which aligns the model’s reasoning capabilities with specific task demands. Second, our results suggest that the most promising application of TTS in MT is not in direct, single-pass translation but in multi-stage workflows, such as a rapid initial draft followed by a more deliberate, computationally-intensive self-correction phase. Future work could explore more dynamic budget allocation strategies and extend to hybrid TTS approaches integrated with external tools like retrieval-augmented generation.
Limitations
While our study provides a comprehensive analysis of test-time scaling in machine translation, we acknowledge several limitations that frame the scope of our conclusions and suggest avenues for future research.
First, our investigation, while encompassing 12 different models, is primarily focused on open-source RM families and a single, smaller proprietary model. The performance characteristics and scaling behaviors of the largest, state-of-the-art proprietary models (e.g., Gemini-2.5-Pro) may differ from our observations. Furthermore, the linguistic diversity of our benchmarks is largely centered around English or Chinese as either a source or target language. Consequently, our findings on the effectiveness of TTS, particularly the interplay with fine-tuning, may not generalize directly to low-resource languages where the reasoning challenges could be substantially different.
Second, our evaluation methodology relies exclusively on automatic and LLM-based metrics. Although we employed a suite of reference-based, reference-free, and specialized LLM-judge metrics to ensure robustness, this approach lacks the nuance of human evaluation. A human assessment would be invaluable for validating our findings, especially on the literary and cultural benchmarks where subtle aspects of style, tone, and appropriateness are critical and may not be fully captured by our current metrics. The potential biases inherent in LLM-as-a-judge frameworks also represent a confounding factor.
Finally, our study implements test-time scaling through a specific budget-forcing mechanism and a simple “wait” token intervention for extrapolation. Other methods for encouraging or extending deliberation, such as alternative prompting strategies or more complex reasoning frameworks, were not explored and could yield different outcomes. Additionally, our analysis is primarily quantitative; we did not perform a qualitative analysis of the content within the models’ reasoning chains. A deeper investigation into what the models are “thinking” could provide valuable insights into why performance plateaus for general-purpose models or why forced extrapolation leads to degradation.
Appendix A Evaluation Prompts
GRB Prompt:
Score the following translation from {src_lang} to {tgt_lang} with respect to the human reference on a continuous scale from 0 to 100, where a score of zero means “no meaning preserved” and score of one hundred means “perfect preservation of meaning, with faithfulness, expressiveness, and elegance”. Only output the score number. {src_lang} source: “{src_text}” {tgt_lang} human reference: “{tgt_text}” {tgt_lang} translation: “{hyp_text}” Score:
GRF Prompt:
Score the following translation from {src_lang} to {tgt_lang} on a continuous scale from 0 to 100, where a score of zero means “no meaning preserved” and score of one hundred means “perfect preservation of meaning, with faithfulness, expressiveness, and elegance”. Only output the score number. {src_lang} source: “{src_text}” {tgt_lang} translation: “{hyp_text}” Score:
GEA100 Prompt:
Please evaluate the following {tgt_lang} translation of an {src_lang} text. Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a {tgt_lang} reader. - 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a {tgt_lang} reader to fully grasp the intended message. - 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a {tgt_lang} reader. - 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a {tgt_lang} reader. - 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a {tgt_lang} reader. Please only output the score number.
GEA5 Prompt:
Please evaluate the following {tgt_lang} translation of an {src_lang} text. Rate the translation on a scale of 0 to 5, where: - 1 point: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a {tgt_lang} reader. - 2 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a {tgt_lang} reader to fully grasp the intended message. - 3 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a {tgt_lang} reader. - 4 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a {tgt_lang} reader. - 5 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a {tgt_lang} reader. Please only output the score number.
Appendix B Evaluation Results of General-purpose Models
Tables 36 summarize the average performance of the General-purpose models across all datasets with respect to the COMET, COMETKiwi, GRB, and GRF metrics, respectively.
Model Budget
0 100 200 300 500 1000 2000 Low High
Cogito-3B 0.712 0.671 0.695 0.698 0.704 0.707 0.704
Cogito-8B 0.765 0.762 0.764 0.763 0.762 0.768 0.769
Qwen3-0.6B 0.702 0.694 0.699 0.695 0.695 0.696 0.689
Qwen3-1.7B 0.760 0.757 0.764 0.763 0.764 0.769 0.764
Qwen3-4B 0.789 0.785 0.789 0.790 0.791 0.791 0.787
Qwen3-8B 0.801 0.800 0.799 0.800 0.799 0.798 0.797
Qwen3-14B 0.805 0.807 0.803 0.806 0.805 0.799 0.797
Qwen3-32B 0.804 0.804 0.802 0.803 0.802 0.802 0.799
Grok-3-Mini 0.794 0.795
Table 3: Average COMET scores of general-purpose models across all datasets with varying thinking budgets.
Model Budget
0 100 200 300 500 1000 2000 Low High
Cogito-3B 0.634 0.599 0.619 0.625 0.629 0.631 0.632
Cogito-8B 0.677 0.676 0.681 0.679 0.680 0.682 0.682
Qwen3-0.6B 0.618 0.617 0.616 0.617 0.614 0.615 0.609
Qwen3-1.7B 0.675 0.671 0.678 0.678 0.679 0.685 0.676
Qwen3-4B 0.706 0.702 0.705 0.707 0.709 0.709 0.700
Qwen3-8B 0.719 0.718 0.717 0.718 0.717 0.718 0.712
Qwen3-14B 0.725 0.724 0.723 0.725 0.725 0.718 0.713
Qwen3-32B 0.724 0.720 0.721 0.722 0.724 0.723 0.717
Grok-3-Mini 0.701 0.701
Table 4: Average COMETKiwi scores of general-purpose models across all datasets with varying thinking budgets.
Model Budget
0 100 200 300 500 1000 2000 low high
Cogito-3B 82.546 78.016 79.590 80.571 81.176 81.497 81.937
Cogito-8B 88.177 88.552 88.312 88.402 87.948 88.432 88.351
Qwen3-0.6B 58.165 57.507 57.518 57.813 57.317 57.372 57.734
Qwen3-1.7B 74.422 74.634 75.522 74.883 75.076 75.678 75.753
Qwen3-4B 84.909 84.921 85.361 85.809 86.165 86.379 86.148
Qwen3-8B 89.204 89.394 89.441 89.623 89.651 89.541 90.001
Qwen3-14B 90.899 90.875 90.707 91.075 91.163 91.253 91.294
Qwen3-32B 90.949 91.103 91.204 91.333 91.340 91.556 91.455
Grok-3-Mini 92.529 92.451
Table 5: Average GRB scores of general-purpose models across all datasets with varying thinking budgets.
Model Budget
0 100 200 300 500 1000 2000 low high
Cogito-3B 80.928 77.468 80.178 81.680 81.577 81.843 80.632
Cogito-8B 89.470 89.550 88.707 88.942 88.082 89.067 89.284
Qwen3-0.6B 56.858 56.943 57.549 57.137 57.232 56.890 56.990
Qwen3-1.7B 74.954 75.377 76.210 75.829 75.723 76.770 76.591
Qwen3-4B 86.238 86.229 86.722 86.965 87.603 87.806 87.577
Qwen3-8B 90.321 90.625 90.648 90.816 91.024 91.123 91.258
Qwen3-14B 92.084 92.016 91.932 92.359 92.503 92.492 92.547
Qwen3-32B 92.122 92.268 92.404 92.608 92.702 92.835 92.782
Grok-3-Mini 93.494 93.594
Table 6: Average GRF scores of general-purpose models across all datasets with varying thinking budgets.
Appendix C Evaluation Results and Thinking Token Statistics of DRT Models on Literary Translation Tasks
Tables 710 present the performance of the DRT models across three literary translation benchmarks with respect to the GEA100, GEA5, GRB, and GRF metrics, respectively. Table 11 shows the token statistics under different budgets. Figure 6 visualizes DRT models’ translation quality (GRB&GRF score, right y-axis) and the actual number of thinking tokens (left y-axis). On the in-domain MetaphorTrans task, there is a positive growth of the quality scores as the thinking budget increases to around 300 tokens, then the score generally stabilizes. While on the other two out-of-domain tasks, the performance fluctuates and shows an overall downward trend as thinking tokens increase.
Figure 6: Performance and actual generated thinking tokens of DRT models across 3 literary translation tasks.
Task Model Budget
100 200 300 400 500 1000 2000
MetaphorTrans DRT-7B 75.08 79.63 80.52 80.72 80.81 80.74 80.89
DRT-8B 71.81 75.36 76.57 77.48 77.70 77.89 78.38
DRT-14B 76.88 81.02 81.62 81.90 82.16 82.39 82.45
LitEval-Corpus DRT-7B 65.09 63.66 65.07 64.38 62.75 68.08 64.47
DRT-8B 58.11 56.92 55.00 55.39 56.10 57.98 58.26
DRT-14B 67.94 69.90 68.97 73.83 74.65 76.07 73.57
WMT24-Literary DRT-7B 79.45 75.48 74.50 77.35 78.03 77.15 76.62
DRT-8B 63.00 59.66 61.57 59.75 67.05 61.18 57.16
DRT-14B 75.34 74.38 77.18 74.06 75.54 81.37 80.01
Table 7: GEA100 scores of DRT models on literary translation tasks with varying thinking budgets.
Task Model Budget
100 200 300 400 500 1000 2000
MetaphorTrans DRT-7B 3.75 3.92 3.95 3.96 3.96 3.96 3.95
DRT-8B 3.58 3.75 3.81 3.85 3.86 3.87 3.88
DRT-14B 3.83 3.97 3.98 4.00 4.00 4.00 4.01
LitEval-Corpus DRT-7B 3.27 3.26 3.27 3.31 3.12 3.29 3.23
DRT-8B 3.02 2.95 2.85 2.96 2.96 3.08 3.01
DRT-14B 3.42 3.45 3.37 3.43 3.58 3.55 3.56
WMT24-Literary DRT-7B 4.02 3.85 3.65 3.58 3.83 3.83 3.64
DRT-8B 2.67 2.93 2.90 3.13 3.09 3.01 2.60
DRT-14B 3.73 3.73 3.58 3.79 3.57 3.86 3.94
Table 8: GEA5 scores of DRT models on literary translation tasks with varying thinking budgets.
Task Model Budget
100 200 300 400 500 1000 2000
MetaphorTrans DRT-7B 92.51 93.01 93.04 92.81 92.73 92.66 92.69
DRT-8B 91.30 91.87 92.07 92.08 92.04 91.85 91.95
DRT-14B 92.36 93.45 93.44 93.25 93.34 93.20 93.27
LitEval-Corpus DRT-7B 67.81 65.00 63.08 63.73 63.02 61.25 59.89
DRT-8B 68.13 63.33 60.89 65.02 61.95 62.55 62.12
DRT-14B 78.52 76.99 74.30 73.73 74.11 74.75 71.26
WMT24-Literary DRT-7B 89.54 88.13 89.06 89.26 88.46 88.25 88.03
DRT-8B 88.83 90.59 87.96 87.10 86.26 85.18 88.76
DRT-14B 90.56 89.81 87.06 88.87 88.53 89.19 89.63
Table 9: GRB scores of DRT models on literary translation tasks with varying thinking budgets.
Task Model Budget
100 200 300 400 500 1000 2000
MetaphorTrans DRT-7B 89.91 90.81 90.84 90.66 90.71 90.66 90.66
DRT-8B 88.25 89.14 89.26 89.37 89.52 89.42 89.65
DRT-14B 89.67 91.35 91.21 91.09 91.44 91.34 91.23
LitEval-Corpus DRT-7B 68.17 65.50 65.32 65.66 64.52 63.25 61.53
DRT-8B 66.01 62.22 60.67 61.75 61.15 62.80 61.43
DRT-14B 75.18 75.88 73.99 72.16 74.13 73.36 71.27
WMT24-Literary DRT-7B 76.06 64.40 70.39 60.50 69.66 60.51 69.00
DRT-8B 39.38 29.54 40.48 40.19 41.83 38.72 35.03
DRT-14B 57.46 60.20 59.01 64.58 74.38 68.64 79.16
Table 10: GRF scores of DRT models on literary translation tasks with varying thinking budgets.
Task Model Budget
100 200 300 400 500 1000 2000
MetaphorTrans DRT-7B 96.50 195.74 295.37 387.48 443.83 476.93 476.94
DRT-8B 95.50 194.87 294.83 390.00 459.57 516.67 514.38
DRT-14B 94.27 195.28 293.20 382.22 439.87 474.03 472.84
LitEval-Corpus DRT-7B 95.97 193.87 293.37 395.45 492.04 831.38 913.88
DRT-8B 91.12 189.58 290.99 393.68 486.73 876.08 1049.21
DRT-14B 94.90 194.57 294.72 391.62 486.15 812.47 919.86
WMT24-Literary DRT-7B 95.89 195.91 290.76 394.55 493.93 966.17 1870.02
DRT-8B 90.71 196.56 268.99 353.15 458.29 902.59 1764.82
DRT-14B 88.43 180.07 258.03 353.63 453.92 876.82 1727.59
Table 11: Actual thinking tokens of DRT models on literary translation tasks with varying thinking budgets.
Appendix D Post-editing Prompts and Detailed Results
D.1 Prompts
We experiment with two variants of post-editing prompts: with and without an additional quality score (QS). Examples are provided below.
Post-editing with QS:
You are a professional translator, and your task is to refine the {tgt_lang} draft translation below based on the {src_lang} source text and its quality evaluation. Please only provide me with the refined translation, without any additional explanations. Source Text: {src_text} Draft Translation: {hyp_text} Quality Score: {quality_score}/100
Post-editing without QS:
You are a professional translator, and your task is to refine the {tgt_lang} draft translation below based on the {src_lang} source text. Please only provide me with the refined translation, without any additional explanations. Source Text: {src_text} Draft Translation: {hyp_text}
D.2 Detailed Results
Tables 12 and 13 report the GRB and GRF scores respectively.
Model Original No QS QS
Budget=0 Budget=500 Budget=1000 Budget=0 Budget=500 Budget=1000
Qwen3-0.6B 69.293 69.756 (+0.463) 69.417 (+0.124) 69.445 (+0.152) 69.112 (-0.181) 69.538 (+0.246) 69.410 (+0.117)
Qwen3-1.7B 82.327 80.913 (-1.415) 82.896 (+0.568) 82.930 (+0.602) 81.438 (-0.889) 82.878 (+0.551) 83.275 (+0.947)
Qwen3-4B 88.695 88.623 (-0.072) 89.734 (+1.039) 89.888 (+1.193) 88.883 (+0.188) 89.725 (+1.030) 89.835 (+1.140)
Qwen3-8B 91.253 91.275 (+0.022) 91.618 (+0.365) 91.747 (+0.494) 91.173 (-0.080) 91.653 (+0.400) 91.657 (+0.404)
Qwen3-14B 92.199 92.241 (+0.042) 92.482 (+0.283) 92.529 (+0.330) 92.294 (+0.095) 92.448 (+0.249) 92.472 (+0.272)
Qwen3-32B 92.197 92.415 (+0.219) 92.502 (+0.305) 92.497 (+0.301) 92.365 (+0.169) 92.282 (+0.085) 92.334 (+0.138)
Table 12: Post-editing GRB score with and without QS at different budgets.
Model Original No QS QS
Budget=0 Budget=500 Budget=1000 Budget=0 Budget=500 Budget=1000
Qwen3-0.6B 67.477 67.757 (+0.281) 67.936 (+0.459) 67.989 (+0.512) 67.987 (+0.510) 68.085 (+0.608) 68.057 (+0.580)
Qwen3-1.7B 81.727 79.159 (-2.568) 81.988 (+0.261) 82.305 (+0.578) 80.260 (-1.467) 82.445 (+0.718) 82.730 (+1.003)
Qwen3-4B 88.588 88.625 (+0.038) 89.936 (+1.349) 90.057 (+1.470) 88.755 (+0.168) 89.885 (+1.298) 90.031 (+1.444)
Qwen3-8B 91.208 90.440 (-0.768) 92.007 (+0.799) 92.018 (+0.810) 89.499 (-1.709) 92.051 (+0.843) 91.933 (+0.725)
Qwen3-14B 92.261 92.613 (+0.352) 92.763 (+0.501) 92.853 (+0.592) 92.642 (+0.381) 92.757 (+0.495) 92.810 (+0.549)
Qwen3-32B 92.256 92.767 (+0.511) 92.923 (+0.667) 92.864 (+0.608) 92.816 (+0.560) 92.768 (+0.512) 92.899 (+0.643)
Table 13: Post-editing GRF score with and without QS at different budgets.

### v2

@@
 Test-Time Scaling of Reasoning Models for Machine Translation
-Zihao Li,1 Shaoxiong Ji,2,3 Jörg Tiedemann1 1University of Helsinki  2 University of Turku  3 ELLIS Institute Finland firstname.lastname@{1helsinki.fi, 2utu.fi}
+Zihao Li,1 Shaoxiong Ji,2,3,1 Jörg Tiedemann1 1University of Helsinki  2 ELLIS Institute Finland  3 University of Turku firstname.lastname@{1helsinki.fi, 3utu.fi}
 1 Introduction
@@
 All these highlight TTS’s value in MT lies in task-specialized models and multi-step self-correction, rather than as a robust strategy for enhancing single-pass translation with general-purpose RMs.
-Large language models (LLMs) have dramatically advanced machine translation (MT), evolving from statistical and neural paradigms to systems capable of handling diverse languages, domains, and complexities with unprecedented accuracy (Lyu et al., 2023; Kocmi et al., 2024; Zhu et al., 2024; Cui et al., 2025).
-Recent developments in Reasoning Models (RMs)—models designed to incorporate structured reasoning processes like Chain-of-Thought (CoT)—have further transformed MT by reframing it as a cognitive task requiring contextual analysis, cultural adaptation, and self-reflection (Liu et al., 2025).
+Large language models (LLMs) have dramatically advanced machine translation (MT), evolving from statistical and neural paradigms to systems capable of handling diverse languages, domains, and complexities with unprecedented accuracy (Lyu et al., 2023; Kocmi et al., 2024; Zhu et al., 2024; Cui et al., 2025; Hendy et al., 2023).
+Recent developments in Reasoning Models (RMs), models designed to incorporate structured reasoning processes like Chain-of-Thought (CoT), have further transformed MT by reframing it as a cognitive task requiring contextual analysis, cultural adaptation, and self-reflection (Liu et al., 2025).
 For instance, RMs can resolve ambiguities in stylized texts and maintain coherence across documents, thereby outperforming traditional LLMs in semantically demanding scenarios (Ye et al., 2025).
@@
 In post-editing (PE) contexts, where models refine their own drafts, TTS might unlock iterative improvements, though this demands rigorous testing across varied benchmarks.
-This paper investigates these open questions through three research questions:
+This paper investigates these open questions by distinguishing between two TTS workflows: Direct Translation (single-pass CoT scaling) and Post-Editing (compute-scaled self-correction).
+We structure our investigation through three research questions:
   • RQ1: How effective is test-time scaling for MT? We examine whether increased inference computation reliably boosts translation quality across general-purpose and fine-tuned MT-specific RMs.
@@
 We investigate if overriding models’ natural stopping points, which further scales up the inference computation, enhances or hinders performance.
-  • RQ3: Does test-time scaling work in post-editing? We evaluate TTS in self-correction scenarios, assessing its role in refining initial translations.
+  • RQ3: Does test-time scaling work in post-editing?
+We evaluate TTS in self-correction scenarios, assessing its role in refining initial translations when being allocated with specific compute budgets.
 To address these questions, we assemble a comprehensive array of MT benchmarks encompassing literary, biomedical, cultural, commonsense, constrained terminology, and retrieval-augmented domains.
@@
 For DRT models fine-tuned on specific domain data, performance improves with budget on in-domain tasks and saturates once models naturally stop increasing their internal token usage, suggesting an emergent alignment between optimal reasoning depth and task demands.
-This alignment largely disappears out of domain.
+This alignment largely disappears out of the domain.
   • We find that forcing a model to reason by inserting a single “wait” beyond its natural stopping point consistently degrades translation quality, highlighting the importance of the model’s intrinsic deliberation process.
@@
   • Cogito: Two models, sized 3B and 8B, were included (Deep Cogito, 2025).
-Cogito-3B and Cogito-8B are trained on top of Llama-3.2-3B and Llama-3.1-8B (Dubey et al., 2024), respectively, and similarly implement hybrid reasoning capabilities with controllable switching between generation and reasoning modes.
+Cogito-3B and Cogito-8B are trained on top of Llama-3.2-3B and Llama-3.1-8B (Grattafiori et al., 2024), respectively, and similarly implement hybrid reasoning capabilities with controllable switching between generation and reasoning modes.
   • DRT: Three models from this family were evaluated.
 These models are fine-tuned from existing LLMs using the training set of MetaphorTrans (Wang et al., 2025a).
-Specifically, DRT-7B, DRT-8B, and DRT-14B are built upon Qwen2.5-7B-Instruct (Qwen et al., 2024), Llama-3.1-8B-Instruct, and Qwen2.5-14B-Instruct respectively.
+Specifically, DRT-7B, DRT-8B, and DRT-14B are built upon Qwen2.5-7B-Instruct (Yang et al., 2024), Llama-3.1-8B-Instruct, and Qwen2.5-14B-Instruct respectively.
 In addition to the open-source models, we included the proprietary model Grok-3-Mini.
@@
 This provides the model with an explicit signal about the quality of the translation it needs to correct, potentially guiding a more targeted reasoning process.
-Figure 2: Average GRB scores of Qwen-3 and Cogito models across all datasets with varying thinking budgets.
+Figure 2: Average GRB scores and average actual thinking tokens of Qwen-3 and Cogito models across all datasets with varying thinking budgets.
 3 Results and Analysis
@@
 These models were evaluated “out-of-the-box" without any fine-tuning on our benchmark datasets.
-Figure 2 plots the average GRB scores across all datasets against varying thinking-token budgets for the Qwen-3 and Cogito model series.
+Figure 2 plots the average GRB scores (right axis) alongside the actual thinking tokens generated (left axis) for the Qwen-3 and Cogito model series.
 After a small initial performance gain when moving from a zero budget to a minimal budget (e.g., 100 tokens), the models’ performance curves almost completely plateau in most cases.
-Further increases in the budget up to 2000 tokens yield no meaningful improvements, indicating that simply allocating more computational steps does not enable the models to produce more refined or accurate translations when they lack specific task-related knowledge.
+A critical observation emerges from the thinking-token curves: despite being allocated budgets up to 2000 tokens, most general-purpose models fail to utilize this capacity.
+The actual reasoning length typically saturates around 600 tokens.
+This suggests that these models reach a “reasoning ceiling” where they autonomously terminate the deliberation process, indicating that simply allocating more computational steps does not enable the models to produce more refined or accurate translations when they lack specific task-related knowledge.
 This conclusion is further corroborated and nuanced by our analysis of the Grok-3-Mini model.
@@
 On the in-domain MetaphorTrans task, we observe a clear and consistent positive correlation between the thinking budget and translation performance.
-As the thinking budget increases from 100 to around 500 tokens, both the number of generated tokens and the GEA scores steadily rise.
+We consider the model’s natural stopping point, where the ‘Actual tokens’ curve plateaus around 500 tokens, as the realistic baseline.
+As the thinking budget increases from 100 to this limit, both the number of generated tokens and the GEA scores steadily rise.
+This monotonic improvement indicates that the model is performing valid, necessary reasoning steps.
 However, beyond a 500-token budget, a critical pattern emerges: the models stop generating more thinking tokens, and concurrently, their performance plateaus.
@@
 Here, the actual thinking tokens continue to scale almost linearly with the budget, indicating the models are using the provided extra capacity to “think" longer.
+[^1] We attribute this distinct behavior to text granularity: unlike the out-of-domain paragraph-level tasks (LitEval-Corpus) where limited context leads models to exhaust reasoning paths and saturate early, the extensive context in document-level translation allows models to continuously expand their reasoning loops to consume the available budget.
 Yet, this extended reasoning does not translate into better performance; the GEA scores remain erratic and show no consistent improvement.
@@
 Machine Translation with Large Language Models
-LLMs have rapidly advanced machine translation, with growing emphasis on open-weight systems that balance translation quality, multilingual coverage, and general-purpose capabilities.
+The application of LLMs to machine translation has witnessed a paradigm shift.
+Foundational research demonstrated that general-purpose models, such as GPT-3, possess remarkable few-shot translation capabilities, challenging traditional supervised systems (Brown et al., 2020).
+Subsequent empirical studies systematically evaluated these capabilities, revealing that while LLMs excel in high-resource languages and specific domains, they often require careful adaptation to match state-of-the-art baselines (Hendy et al., 2023; Zhu et al., 2024).
+To address these limitations, researchers have developed sophisticated in-context learning strategies (Agrawal et al., 2023) and diverse prompting techniques (Vilar et al., 2023).
+Notably, recent work by Wu et al.
+(2025) highlights the efficacy of simple re-translation prompts over complex reasoning in general LLMs, suggesting that iterative refinement can significantly boost performance without elaborate chains of thought.
+Concurrently, the field is moving towards open-weight models that balance broad multilingual competence with translation specialization.
 Seed-X (Cheng et al., 2025) introduces a 7B-parameter open-source family of translation-oriented LLMs trained on large-scale monolingual and bilingual corpora across 28 languages, achieving performance competitive with closed-source systems such as GPT-4o (Hurst et al., 2024) and Gemini-2.5 (Comanici et al., 2025).
-Similarly, Hunyuan-MT (Zheng et al., 2025) develops 7B-parameter models: Hunyuan-MT and Hunyuan-MT-Chimera, the latter integrates multiple outputs under a “slow thinking” paradigm to yield higher-quality translations, and ranks first in the WMT2025 shared task across 30 of 31 directions.
+Similarly, Hunyuan-MT (Zheng et al., 2025) develops 7B-parameter models: Hunyuan-MT and Hunyuan-MT-Chimera, the latter integrates multiple outputs under a “slow thinking” paradigm to yield higher-quality translations, and ranks first in the WMT2025 shared task across 30 of 31 directions (Kocmi et al., 2025).
 Tower+ (Rei et al., 2025) addresses the trade-off between translation specialization and general-purpose ability by combining continued pretraining, supervised fine-tuning, preference optimization, and reinforcement learning with verifiable rewards.
@@
 A deeper investigation into what the models are “thinking” could provide valuable insights into why performance plateaus for general-purpose models or why forced extrapolation leads to degradation.
+Acknowledgments
+The work has received funding from the Digital Europe Programme under grant agreement No 101195233 (OpenEuroLLM).
+The authors wish to acknowledge CSC - IT Center for Science, Finland, and LUMI supercomputers, owned by the EuroHPC Joint Undertaking, for providing computational resources.
+Shaoxiong Ji received support from the CA21167 COST action UniDive, funded by COST (European Cooperation in Science and Technology).
 Appendix A Evaluation Prompts